The function ∫ is defined by
∫(x)=

-6+x
______
7+3x
Find ∫(y+2).
∫(y+2)=???
I thought the answer might be something like...
∫(y+2)=
y-4
______
3y+13

Answer by rapture(86)

f(x)= (-6 + x)/(7 + 3x)


Find f(y+2).


Replace every x you see with (y + 2).


f(y + 2) = (-6 + (y + 2))/(7 + 3(y + 2))


f(y + 2) = (y - 4)/(7 + 3y + 6)


Answer: f(y + 2) = (y - 4)/(3y + 13)


Basically,the question is asking you to evaluate the original function. This is a very important skill to master when taking math courses.
harpazo1965@gmail.com